Cash Poor, Inc. (CPI) buys on terms of 2/10, net 45. CPI does not take the discount, and it generally pays after 50 days. What is the nominal annual percentage cost of its non–free trade credit, based on a 365–day year? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.
Answer #1

Discount Rate = 2%
Discount Period = 10 days
Credit Period = 50 days

Nominal Cost of Credit = [Discount Rate / (1 - Discount Rate)] * [365 / (Credit Period - Discount Period)]
Nominal Cost of Credit = [0.02 / (1 - 0.02)] * [365 / (50 - 10)]
Nominal Cost of Credit = [0.02 / 0.98] * [365 / 40]
Nominal Cost of Credit = 0.020408 * 9.1250
Nominal Cost of Credit = 0.1862 or 18.62%

So, nominal annual percentage cost of trade credit is 18.62%
Answer #2

To calculate the nominal annual percentage cost of non-free trade credit, we need to consider the cost of not taking the discount and paying after 50 days instead of the discount period.

Given: Terms: 2/10, net 45 CPI pays after 50 days

To calculate the nominal annual percentage cost, we can use the following formula:

Nominal Annual Percentage Cost = (Discount % / (1 - Discount %)) * (365 / (Payment Days - Discount Days))

Where: Discount % = 2/100 = 0.02 (since the terms are 2/10) Payment Days = 50 days Discount Days = 10 days

Now, plug in the values:

Nominal Annual Percentage Cost = (0.02 / (1 - 0.02)) * (365 / (50 - 10)) Nominal Annual Percentage Cost = (0.02 / 0.98) * (365 / 40) Nominal Annual Percentage Cost = 0.020408 * 9.125 Nominal Annual Percentage Cost = 0.186224

Round the answer to two decimal places:

Nominal Annual Percentage Cost = 18.62%

So, the nominal annual percentage cost of non-free trade credit for Cash Poor, Inc. is 18.62%.


answered by: mervetokaz
Answer #3

To calculate the nominal annual percentage cost of non-free trade credit, we need to find the equivalent interest rate for the credit terms. The credit terms are 2/10, net 45, which means that a 2% discount is offered if payment is made within 10 days, otherwise, the full payment is due within 45 days.

Since CPI does not take the discount and generally pays after 50 days, it effectively borrows money for an additional 5 days beyond the net 45 terms.

To calculate the nominal annual percentage cost, we can use the formula:

Nominal Annual Percentage Cost = (Discount % / (1 - Discount %)) * (365 / (Number of Days credit is extended - Discount Period))

Where: Discount % = 2% (from the credit terms) Number of Days credit is extended = 50 days (CPI pays after 50 days) Discount Period = 10 days (time to take the discount)

Now, let's calculate the nominal annual percentage cost:

Discount % = 2% = 0.02 Number of Days credit is extended = 50 days Discount Period = 10 days

Nominal Annual Percentage Cost = (0.02 / (1 - 0.02)) * (365 / (50 - 10)) Nominal Annual Percentage Cost ≈ 0.0611

Therefore, the nominal annual percentage cost of non-free trade credit for CPI is approximately 6.11%.

answered by: Hydra Master